# Ultrafast Sampling-based Kinodynamic Planning via Differential Flatness

### 17 Mar 2026 v1

@@ VI Evaluation
Figure 4: Visualization of our trajectories with DynoBench [ortizharo2025iDbAstar] benchmarking problems: the trajectories from our kinodynamic RRT* and our version of SST* are plotted as blue and magenta solid curves, respectively, while the iDb-A* and SST* baselines’ are shown as orange dashed and green dotted curves, respectively.
@@
Robot Problem Our RRT-Connect Our SST* iDb-A* SST*
Length Planning Time Length Planning Time Length Planning Time Length Planning Time
(m) (ms) (m) (ms) (m) (ms) (m) (ms)
Unicycle Bugtrap $12.39$ $\boldsymbol{6.0}$} $12.51$ $48.6$ $\boldsymbol{11.51}$} $400$ $12.38$ $200$
Wall $4.82$ $\boldsymbol{0.6}$} $5.24$ $11.0$ $\boldsymbol{3.89}$} $100$ $4.97$ $100$
2D Quadrotor Bugtrap $11.46$ $\boldsymbol{2.7}$} $11.67$ $45$ $9.69$ $10800$ $\boldsymbol{9.64}$} $28700$
Hole $5.16$ $\boldsymbol{0.4}$} $4.93$ $270$ $4.62$ $400$ $\boldsymbol{4.59}$} $17700$
3D Quadrotor Block $8.08$ $\boldsymbol{0.9}$} $8.61$ $1.4$ $\boldsymbol{7.94}$} $4600$ $9.32$ $45800$
Window $5.49$ $\boldsymbol{0.5}$} $6.28$ $22.6$ $\boldsymbol{5.12}$} $2700$ $6.72$ $46300$
@@
TABLE II: Planning performance with a Franka Emika Panda robot over $100$ runs of our kinodynamic RRT-Connect with different environments in MotionBenchMaker [chamzas2022-motion-bench-maker]. For a baseline, we use a collision-free geometric path, provided by a SIMD-parallelized geometric RRT-Connect [thomason2024vamp], followed by time parameterization (TOPP-RA [pham2018toppra]). Overall, our kinodynamic RRT-Connect generates a dynamically feasible trajectory faster than the baseline in almost all environments. Our trajectory is slightly longer but verified to be collision-free while the baseline’s trajectory is colliding in approximately $30\%$ of the cases, due to the time parameterization. Better metrics are shown in bold, while the collision risk is marked as red if positive.
Env. Metrics Our kinodynamic RRT-Connect (AkinoPDF) Geometric RRT-Connect (VAMP) + TOPP-RA
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.34$} $0.48$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.32$} $0.47$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.29$} $0.46$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.13$} $0.34$ 0 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.40$} $0.49$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.42$} $0.49$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
    $\triangleright$ RRT-Connect time $3.45$ $18.00$ $0.08$ $0.2$ $0.70$ $14.54$ $\boldsymbol{0.18}$} $0.36$ $\boldsymbol{0.05}$} $\boldsymbol{0.07}$} $\boldsymbol{0.15}$} $\boldsymbol{0.83}$}
    $\triangleright$ Simpl./TOPP-RA time $\boldsymbol{0.04}$} $0.03$ $\boldsymbol{0.02}$} $\boldsymbol{0.04}$} $\boldsymbol{0.05}$} $\boldsymbol{0.09}$} $4.24$ $1.22$ $3.28$ $4.02$ $4.99$ $6.58$
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0.33$} $0.47$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
Figure 5: Visualization of trajectories generated by our kinodynamic RRT-Connect (bottom) and by the baseline planner, RRT-Connect + TOPP-RA (top), for a Franka Emika Panda robot in: (a) cage, (b) box, (c) bookshelf thin, and (d) table pick environments. For each environment, the top image shows a trajectory from the baseline with collided configurations in red, while the bottom image demonstrates that our trajectory is collision-free.
In this section, we will evaluate the effectiveness of our AkinoPDF approach on different robot platforms, from low to high dimensional state spaces in both simulation and real experiments. The results from our planners and the baselines are both reported from an Intel i7-6900K CPU for a fair comparison. As our method focuses on fast kinodynamic planning on CPUs, comparison to other GPU-based motion planners is out of scope of this paper.
VI-A Kinodynamic planning for low-d o f systems
[…]
For each platform, we consider $2$ DynoBench environments, shown in Fig. 4: Bugtrap and Wall for the unicycle, Bugtrap and Hole for the 22D quadrotor, and Window and Obstacle for the 33D quadrotor. To be compatible with simd-based collision checking, e.g.[thomason2024vamp], the obstacles in each environment are represented by a set of spheres, while the unicycle and quadrotor’s geometries are both modeled as single spheres.
As we obtain closed-form solutions for both BVPs and dynamics propagation for tree expansion, we implement two planners: a kinodynamic version of RRT-Connect [kuffner2000rrtconnect] with BVP solutions in Sec. V-C1 and an SST* [li2016sst] with dynamics propagation in Sec. V-C2. Our kinodynamic RRT-Connect replaces a line segment connecting two nodes in an RRT tree by our closed-form BVP solution (local “flat” path) in the FlatExtend subroutine (Alg. 2), using Eq. 24 with an optimal $T$ from Eq. 25 and $\rho=1$. In our SST* version, the FlatExtend subroutine uses Eq. 28 to propagate robot dynamics in closed forms. While the original SST* keeps expanding the tree using dynamics propagation until it reaches the goal, our version takes advantage of our minimum-time local “flat” path in (24) to check whether there is a direct connection from an existing node to the goal. The resulting trajectories are simplified by skipping unnecessary local paths using Alg. 4. We compare our approach with two baseline planners: iDb-A* [ortizharo2025iDbAstar] and SST* [li2016sst] from OMPL [sucan2012the-open-motion-planning-library], provided by the DynoBench benchmark [ortizharo2025iDbAstar]. We maintain the aforementioned robot state and control constraints in our planners, enforced via our parallelized collision checking algorithm in Alg. 3, and in the baselines for a fair comparison.
As each method has their own definition of trajectory costs, we choose the trajectory length and planning time as common metrics for comparison, shown in Table I. Qualitatively, Fig. 4 plots the trajectories generated by our RRT-Connect (blue solid curves), our SST* (magenta solid curves), iDb-A* (orange dashed curves), and SST* (green dotted curves). The baselines iDb-A* and SST* generate shorter trajectories but often take several seconds to plan, especially for complicated robot dynamics such as quadrotors’. For all $3$ robot platforms, our RRT-Connect shows significantly faster planning times compared to the baselines, in just a few milliseconds, offering the capability of real-time kinodynamic planning. Our SST* planner is slightly slower than our RRT-Connect since it is restricted to a constant control input for each local “flat” path instead of directly calculating a time-parameterized optimal control value. However, it is still significantly faster than the baselines in most cases. Meanwhile, our methods maintain comparable trajectory lengths, within 10–20% of the shortest length across the benchmarking problems. While our generated trajectories can be further optimized by combining with other optimization-based motion planners, e.g., as initial solutions, this is out of the scope of our paper and left for future work.
VI-B Kinodynamic planning for high-d o f manipulators
In our experiments, we use our kinodynamic RRT-Connect planner, described in Sec. VI-A with trajectory simplification in Alg. 4. Our baseline is a common approach that generates a time-parameterized trajectory, e.g., using TOPP-RA [pham2018toppra], from a collision-free geometric path, under kinematic constraints such as velocity and acceleration limits. The geometric path is provided by a simd-parallelized geometric RRT-Connect planner from VAMP [thomason2024vamp]. For each environment, we use both our planner and the baseline to solve the pre-generated benchmarking problems and report the results in Table II. Our planner’s total planning time is calculated as the sum of the RRT-Connect time and trajectory simplification time. Meanwhile, the baseline’s total planning time includes the RRT-Connect time, path simplification time, and TOPP-RA time. As our trajectory and the baseline’s are optimized for different cost functions, we use their length as a common metrics, similar to Sec. VI-A. As TOPP-RA [pham2018toppra] only fits a time-parameterized trajectory to geometric path, i.e., a sequence of waypoints, it does not consider any collision avoidance constraints. While the geometric path is collision-free, there is no guarantee that the time-parameterized trajectory from TOPP-RA will be valid. Therefore, we sample the resulting trajectories, check for collision and compare the collision risk, measured by the percentage of collided trajectories.
Table II compares our kinodynamic planner and the baseline, in terms of the total planning time, the final trajectory’s length, and the collision risk, for each environments and overall. On average, we are able to generate a valid trajectory in around $3.5ms$, and in less than $1ms$ for $75\%$ of the problems, while the baseline’s total planning time is consistently around $4.5ms$ for almost all problems. While VAMP can generate a collision-free geometric path fast, the baseline’s total planning time is dominated by the TOPP-RA time. Our approach, instead, spends most of the time in the kinodynamic RRT-Connect planner, as we directly enforce the dynamics constraints in the RRT tree construction. Across all the problems, the baseline offers shorter trajectory but, more importantly, poses a collision risk of $\sim 30\%$ on average, highlighting the benefit of our approach with guarantees of collision-free trajectories up to the sampling resolution. Fig. 5 visualizes the trajectories from our kinodynamic RRT (bottom) with no collisions and from the baseline (top) with collided configurations shown in red.

### 21 Apr 2026 v2

@@ VII Evaluation
Figure 5: Visualization of our trajectories with DynoBench [ortizharo2025iDbAstar] benchmarking problems: the trajectories from our kinodynamic FLASK-RRTConnect and our BVP-augmented FLASK-SST are plotted as blue and magenta solid curves, respectively, while those from the iDb-A* and SST* baselines are shown as orange dashed and green dotted curves, respectively, and the Kino-PAX trajectories are plotted as purple dash-dotted curves.
In this section, we will evaluate the effectiveness of our FLASK framework, applied on different planners and on different robot platforms, from low to high dimensional state spaces in both simulation and real experiments. As our method focuses on fast kinodynamic planning on CPUs, the results from our planners and the baselines are reported on an Intel i7-6900K CPU for a fair comparison. As a reference, we also provide the results from the GPU-based Kino-PAX [perrault2025kino] planner, which is run on a GeForce RTX 4070 GPU.
VII-A Kinodynamic planning for low-d o f systems
[…]
@@ VII-A Kinodynamic planning for low-d o f systems
Robot Problem FLASK- FLASK-SST FLASK-SST FLASK-RRT iDb-A* Orig. SST* Kino-PAX
RRTConnect (BVP-aug.) (SIMD-only) (DP-based) (DynoBench) (DynoBench) (GPU-based)
Traj. Plan. Traj. Plan. Traj. Plan. Traj. Plan. Traj. Plan. Traj. Plan. Traj. Plan.
Len. Time Len. Time Len. Time Len. Time Len. Time Len. Time Len. Time
(m) (ms) (m) (ms) (m) (ms) (m) (ms) (m) (ms) (m) (ms) (m) (ms)
Uni. Bugtrap $12.39$ $\boldsymbol{6.0}$} $12.51$ $48.6$ 12.27{{12.27}} 407{{407}} 12.25{{12.25}} 565{{565}} $\boldsymbol{11.51}$} $400$ $12.38$ $200$ 13.77{{13.77}} 5.8{{5.8}}
Wall $4.82$ $\boldsymbol{0.6}$} $5.24$ $11.0$ 5.27{{5.27}} 29.5{{29.5}} 5.61{{5.61}} 24.8{{24.8}} $\boldsymbol{3.89}$} $100$ $4.97$ $100$ 5.28{{5.28}} 2.6{{2.6}}
Quad. Bugtrap $11.46$ $\boldsymbol{2.7}$} $11.67$ $45$ 11.46{{11.46}} 388{{388}} 11.50{{11.50}} 909{{909}} $9.69$ $10800$ $\boldsymbol{9.64}$} $28700$ 18.59{{18.59}} 19.7{{19.7}}
(2D) Hole $5.16$ $\boldsymbol{0.4}$} $4.93$ $270$ 5.14{{5.14}} 1813{{1813}} 5.24{{5.24}} 2133{{2133}} $4.62$ $400$ $\boldsymbol{4.59}$} $17700$ 17.82{{17.82}} 124.9{{124.9}}
Quad. Block $8.08$ $\boldsymbol{0.9}$} $8.61$ $1.4$ 9.56{{9.56}} 4752{{4752}} 8.75{{8.75}} 7602{{7602}} $\boldsymbol{7.94}$} $4600$ $9.32$ $45800$ 10.36{{10.36}} 21.4{{21.4}}
(3D) Window $5.49$ $\boldsymbol{0.5}$} $6.28$ $22.6$ 6.96{{6.96}} 629{{629}} 7.98{{7.98}} 2631{{2631}} $\boldsymbol{5.12}$} $2700$ $6.72$ $46300$ 16.41{{16.41}} 178.6{{178.6}}
Figure 6: Visualization of trajectories generated by our FLASK-RRTConnect (top) and by the baseline planner, geometric RRTConnect + TOPP-RA (bottom), for a Franka Emika Panda robot in: (a) cage, (b) box, (c) bookshelf thin, and (d) table pick environments. For each environment, the top image shows a trajectory from the baseline with collided configurations in red, while the bottom image demonstrates that our trajectory is collision-free.
For each platform, we consider $2$ DynoBench environments, shown in Fig. 5: Bugtrap and Wall for the unicycle, Bugtrap and Hole for the 22D quadrotor, and Window and Obstacle for the 33D quadrotor. To be compatible with simd-only collision checking, e.g.[thomason2024vamp], the obstacles in each environment are represented by a set of spheres, while the unicycle and quadrotor’s geometries are both modeled as single spheres.
We apply our FLASK framework to obtain two planners: FLASK-RRTConnect with BVP solutions in Sec. V-C1 and three variants of FLASK-SST with dynamics propagation in Sec. V-C2. Our kinodynamic FLASK-RRTConnect replaces a line segment connecting two nodes in an RRT tree by our closed-form BVP solution (local “flat” path) in the FlaskExtend subroutine (Alg. 2), using Eq. 25 with an optimal $T^{*}$ from Eq. 26 and $\rho=1$. For FLASK-SST, our first variant, BVP-augmented FLASK-SST, uses Eq. 29 to propagate robot dynamics in closed forms and takes advantage of our minimum-time local “flat” path in (25) to check whether there is a direct connection from an existing node to the goal. Meanwhile, our second variant is a SIMD-only FLASK-SST that keeps expanding the tree using vectorized collision checking via SIMD, without leveraging the BVP solutions. To highlight the “wandering” effect in propagation-based planners, our third variant disables best-state selection and tree pruning in SST*, effectively leading to a DP-based FLASK-RRT, where dynamics propagation (DP) with a sampled control is used to extend the RRT tree. The resulting trajectories are simplified by skipping unnecessary local paths using Alg. 4. We compare our approach with three baseline planners: iDb-A* [ortizharo2025iDbAstar], SST* [li2016sst] from OMPL [sucan2012the-open-motion-planning-library], both without parallelized collision checking, and a GPU-based coarse-grained parallelized kinodynamic planner, Kino-PAX [perrault2025kino]. We maintain the aforementioned robot state and control constraints in our planners, enforced via our parallelized collision checking algorithm in Alg. 3, and in the baselines for a fair comparison.
As each method has its own definition of trajectory costs, it is challenging to compare the quality of the trajectories. However, most cost functions describe a combination of trajectory duration, length, control efforts, and even velocity, acceleration, or higher-order terms, all of which intuitively and indirectly prioritize shorter trajectories. Therefore, we choose the trajectory length and planning time as common metrics for comparison, shown in Table I. For anytime planners such as iDb-A* [ortizharo2025iDbAstar], SST* [li2016sst], we report the time of the first solution and the length of the final solution within a $180$-second limit. Qualitatively, Fig. 5 plots the trajectories generated by our FLASK-RRTConnect (blue solid curves), our BVP-augmented FLASK-SST (magenta solid curves), iDb-A* (orange dashed curves), SST* (green dotted curves), and Kino-PAX (dash-dotted purple curves).
VII-A1 Comparison to the baselines
The baselines iDb-A* and SST* generate shorter trajectories but often take several seconds to plan, especially for complicated robot dynamics such as quadrotors. Kino-PAX [perrault2025kino] achieves planning times in the millisecond range by parallelizing the tree expansion on GPUs but returns much longer trajectories, possibly due to the “wandering” effect” from dynamics propagation with sampled suboptimal control. While it achieves similar planning times for simple dynamics such as unicycles, it is $\sim$8–300 times slower than our FLASK-RRTConnect for more complicated systems such as quadrotors. For all $3$ robot platforms, our FLASK-RRTConnect achieves significantly faster planning times than the baselines, in just a few milliseconds, offering the capability of real-time kinodynamic planning. Our BVP-augmented FLASK-SST variant is slightly slower than our FLASK-RRTConnect since it is restricted to a constant control input for each local “flat” path instead of directly calculating a time-parameterized optimal control value. However, it is still significantly faster than the baselines in most cases. Meanwhile, our methods maintain comparable trajectory lengths, within 10–20% of the shortest length across the benchmarking problems. Intuitively, this is because the cost function (20) prioritizes shorter trajectories, which leads to lower total control effort and trajectory time. While our generated trajectories can be further optimized by combining with other optimization-based motion planners, e.g., as initial solutions, this is out of the scope of our paper and left for future work.
@@ VII-A1 Comparison to the baselines
TABLE II: Planning performance with a Franka Emika Panda robot over $100$ runs of our FLASK-RRTConnect with different environments in MotionBenchMaker [chamzas2022-motion-bench-maker]. For a baseline, we use a collision-free geometric path, provided by a SIMD-parallelized geometric RRTConnect [thomason2024vamp], followed by time parameterization (TOPP-RA [pham2018toppra]). Overall, our kinodynamic FLASK-RRTConnect generates a dynamically feasible trajectory faster than the baseline in almost all environments. Our trajectory is slightly longer but verified to be collision-free while the baseline’s trajectory is colliding in approximately $30\%$ of the cases, due to the time parameterization. Better metrics are shown in bold, while the collision risk is marked as red if positive.
Env. Metrics Our kinodynamic FLASK-RRTConnect Geometric RRTConnect (VAMP) + TOPP-RA
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.34 $0.48$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.32 $0.47$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.29 $0.46$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.13 $0.34$ 0 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.40 $0.49$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.42 $0.49$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
    $\triangleright$ RRTConnect time $3.45$ $18.00$ $0.08$ $0.2$ $0.70$ $14.54$ $\boldsymbol{0.18}$} $0.36$ $\boldsymbol{0.05}$} $\boldsymbol{0.07}$} $\boldsymbol{0.15}$} $\boldsymbol{0.83}$}
    $\triangleright$ Post./TOPP-RA time $\boldsymbol{0.04}$} $0.03$ $\boldsymbol{0.02}$} $\boldsymbol{0.04}$} $\boldsymbol{0.05}$} $\boldsymbol{0.09}$} $4.24$ $1.22$ $3.28$ $4.02$ $4.99$ $6.58$
@@
Collision risk $\boldsymbol{0}$} 0 $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} $\boldsymbol{0}$} 0.33 $0.47$ 0 0 $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1$}
@@
VII-A2 The roles of BVP solutions and parallelized collision checking
All of our planners in the experiments use parallelized collision checking but integrate the BVP solution at different levels: (i) our FLASK-RRTConnect uses the BVP solution to find all the local paths; (ii) our BVP-augmented FLASK-SST only uses the BVP solution to find a shortcut to the goal; (iii) our SIMD-only FLASK-SST only leverages the vectorized collision checking without using the BVP solution; and (iv) our DP-based FLASK-RRT simply expands the planning tree using dynamics propagation without best-state selection and pruning benefits from SST*.
Parallelized collision checking clearly improves the planning times. Compared to iDb-A* [ortizharo2025iDbAstar] and the original SST* [li2016sst], checking the local path for collision in parallel reduces the planning times from tens of seconds to a few seconds (see Table I) for robots with complicated dynamics such as quadrotors. This is the case even when the BVP solution is not used as in our SIMD-only FLASK-SST or there is no best-state selection and tree pruning as in our DP-based FLASK-RRT. As propagation-based planners, our SIMD-only FLASK-SST and DP-based FLASK-RRT still wander through the state space, a few small time steps at a time, causing longer planning times and trajectory lengths, as shown in Table I. This effect is even worse in DP-based FLASK-RRT as the best-state selection and pruning techniques used in SST* are not available.
However, our unique integration of BVP solutions and parallelized collision checking is the key to driving the planning times further down to the (sub-)millisecond range. By employing the BVP solution to find a shortcut from a node to the goal, our BVP-augmented FLASK-SST generates a trajectory in just milliseconds, significantly faster than SIMD-only FLASK-SST. This illustrates the potential of our framework to improve existing propagation-based planners by addressing the common “wandering” effects via our BVP solution. Taking a further step, our FLASK-RRTConnect fully integrates the BVP solution to find all local paths during tree expansion, and therefore, leads to less than a millisecond of planning time in many cases. This illustrates our approach’s ability to transform any existing sampling-based geometric planner into an ultrafast kinodynamic planner, requiring only widely available CPUs.
VII-B Kinodynamic planning for high-d o f manipulators
In our experiments, we use our FLASK-RRTConnect planner, described in Sec. VII-A with trajectory simplification in Alg. 4. Our baseline is a common approach that generates a time-parameterized trajectory, e.g., using TOPP-RA [pham2018toppra], from a collision-free geometric path, under kinematic constraints such as velocity and acceleration limits. The geometric path is provided by a simd-parallelized geometric RRTConnect planner from VAMP [thomason2024vamp]. For each environment, we use both our planner and the baseline to solve the pre-generated benchmarking problems and report the results in Table II. Our planner’s total planning time is calculated as the sum of the RRTConnect time and trajectory simplification time. Meanwhile, the baseline’s total planning time includes the RRTConnect time, path simplification time, and TOPP-RA time. As our trajectory and the baseline’s are optimized for different cost functions, we use their length as a common metrics, similar to Sec. VII-A. As TOPP-RA [pham2018toppra] only fits a time-parameterized trajectory to geometric path, i.e., a sequence of waypoints, it does not consider any collision avoidance constraints. While the geometric path is collision-free, there is no guarantee that the time-parameterized trajectory from TOPP-RA will be valid. Therefore, we sample the resulting trajectories, check for collision and compare the collision risk, measured by the percentage of collided trajectories.
Table II compares our kinodynamic planner and the baseline, in terms of the total planning time, the final trajectory’s length, and the collision risk, for each environment and overall. On average, we are able to generate a valid trajectory in around $3.5ms$, and in less than $1ms$ for $75\%$ of the problems, while the baseline’s total planning time is consistently around $4.5ms$ for almost all problems. While VAMP can generate a collision-free geometric path fast, the baseline’s total planning time is dominated by the TOPP-RA time. Our approach, instead, spends most of the time in the FLASK-RRTConnect planner, as we directly enforce the dynamics constraints in the RRT tree construction. Across all the problems, the baseline offers a shorter trajectory but, more importantly, poses a collision risk of $\sim 30\%$ on average, highlighting the benefit of our approach with guarantees of collision-free trajectories up to the sampling resolution. Fig. 6 visualizes the trajectories from our FLASK-RRTConnect (top) with no collisions and from the baseline (bottom) with collided configurations (red).
VII-C Real experiments
Finally, we verify the benefit of our approach for online motion planning on a real Universal Robots (UR5) platform in a cluttered environment. The environment consists of multiple static and dynamic objects, perceived by an Intel Realsense camera with a top-down view. The depth image from the camera is used to generate a set of spheres stored in a collision-affording point tree (CAPT) compatible with simd-parallelized motion planning [ramsey2024-capt]. The spheres represent the obstacles in the environment in our Alg. 3. Our FLASK-RRTConnect planner in Sec. VII-A is used to find a trajectory from the start configuration to the goal.
VII-C1 Kinodynamic Planning versus Geometric Planning
We first consider a “pick and place” scenario in Fig. 1, where the robot generates a motion plan from its start position to a “pick” position near a granola box, and then move to a “place” position above a basket. The environment is cluttered with multiple objects, creating narrow passages such that a minor deviation from the planned trajectory can cause collisions. Therefore, accurate tracking is crucial for the safe realization of this “pick and place” task. The baseline is a geometric path, generated by a SIMD-parallelized RRTConnect planner from VAMP [thomason2024vamp]. The path is sampled with a constant open-loop velocity, chosen such that the task’s duration is similar to ours for a fair comparison. A closed-loop velocity control input is computed by Ruckig [berscheid2021ruckig] and sent to the UR5 robot to execute.
Fig. 7 plots the tracking error during the experiments. Our trajectory is dynamically feasible, allowing the robot to smoothly and accurately execute the task without any collisions (Figs. 7(a) and 8(a)). Meanwhile, the geometric path has sharp turns, leading to overshoots and fluctuations (Fig. 7(b)). As a result, the robot collides with two nearby boxes while trying to turn, as seen in Fig. 8(b). The overshoots can be observed in Fig. 7(c) with two spikes in total tracking error $\|\mathbf{q}_{actual}-\mathbf{q}_{desired}\|$, at times $t\approx 4$s and $t\approx 9$s, where $\mathbf{q}_{actual}$ and $\mathbf{q}_{desired}$ denote the actual and desired values of the joint angles from the motion plans. Meanwhile, our tracking error stays low without any spikes, highlighting the benefits of having a dynamically feasible trajectory from our kinodynamic planner for accurate task executions.